Find 2.48 ÷ 4 Find 2.48 ÷ 4 Find 2.48 ÷ 4 Find 2.48 ÷ 4

Find 
2.48 ÷ 4
Find 
2.48 ÷ 4
Find 
2.48 ÷ 4
Find 
2.48 ÷ 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (248/100) ÷ 4
Then,
= (248/100) × (1/4)
= (62/100) × (1/1)
= (62/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places.
Then,
= 0.62

Thanks